OpenStudy (abbycross167):

Can someone please check my work? Please? I would really appreciate it... Question~ Solve the systems using the elimination method. y = x - 1 y = -x + 7

OpenStudy (abbycross167):

My Answer~ -2x+y=-1 X+y=7 = -2x+y=-1 Multiply by 1 X+y=7 multiply by 2 = 2x-1=-x+7 3x-1=7 3x=8 X=8/3 Y=13/3 Final answer~ (8/3, 13,3)

OpenStudy (abbycross167):

What my teacher said was wrong~ your first equation under Answer doesn't make sense

OpenStudy (kenshin):

how did you get -2x+y=1? there was only 1 x on both equations wasn't there?

OpenStudy (kenshin):

-2x+y=-1 i meant

OpenStudy (abbycross167):

well I wasn't sure how to answer this question, so I asked my brother for help and he told me to start with that... He didn't tell me how he got it

OpenStudy (kenshin):

do you understand how to eliminate x from both equations by adding them together?

OpenStudy (abbycross167):

somewhat...

OpenStudy (kenshin):

ok so we have (1) y = x - 1, and (2) y = - x + 7 what's (3) if you add (1) and (2) together?

OpenStudy (abbycross167):

So 3+1+2=6 is that what you meant to do?

OpenStudy (kenshin):

no those are just names for each equation I could have picked (A) (B) and (C) or (T) (F) (G) or whatever, by (1) I mean't the entire equation y=x-1, by (2) I meant the equation y=-x+7 and (3) is adding (1) and (2) together, do you know how to do that?

OpenStudy (kenshin):



OpenStudy (abbycross167):

I'm not sure... I'm sorry I have a hard time with math

OpenStudy (kenshin):

ok what's y + y? and x + -x? and 1+7?

OpenStudy (abbycross167):

y+y=y or 2y?

OpenStudy (kenshin):

2y, and x+ -x?

OpenStudy (abbycross167):

-2x?

OpenStudy (kenshin):

no, what's 1+ (-1)

OpenStudy (abbycross167):

0

OpenStudy (kenshin):

that's right, so do you get how we can get 2y = 0x + (1+7) now? and what is y now?

OpenStudy (abbycross167):

I don't think I understand what you mean by "what is y now?" wouldn't it be 2y still?

OpenStudy (kenshin):

what's 2y equal to right now then?

OpenStudy (abbycross167):

8y?

OpenStudy (kenshin):

no we had 2y = 0x + 1 + 7, what is 2y alone equal to?

OpenStudy (abbycross167):

2y

OpenStudy (kenshin):

and that is equal to what on the right hand side?

OpenStudy (abbycross167):

0x+1=2? I'm not sure if that is how I do it...